12 games were won out of 16. What percent of the games were won?
My solution:
16-12=4
4/16= 1/4 convert 4/1= .25; therefore they won 25% of their games, but the answer key says that it is 75%. I just want to know if i'm right or wrong; if not please instruct me where the error is; thank you.

Answer by jim_thompson5910(35256)

Percent of games won = (# of games won)/(# of games total) = 12/16 = 0.75 which converts to 75%